The demand for labor in both Occupations A and B is L_D = 15 - 2W, where L_D is the number of workers demanded for that occupation, in thousands, at wage rate W. The supply of labor for Occupation A is L_A = -4 + W and for Occupation B is L_B = -2 + (1.2)W. What is the compensating wage differential between the two occupations?
Answer #1

In occupation A, equilibrium occurs where LD = LA
So, 15 - 2W = -4 + W
So, 2W + W = 15 + 4
So, 3W = 19
So, WA = 19/3 = 6.33

In occupation B, equilibrium occurs where LD = LB
So, 15 - 2W = -2 + 1.2W
So, 2W + 1.2W = 15 + 2
So, 3.2W = 17
So, WB = 17/3.2 = 170/32 = 5.31

Compensating Wage differential = WA - WB = 6.33 - 5.31 = 1.02